### v2

# RELIEF: Reinforcement Learning Empowered Graph Feature Prompt Tuning

Abstract.
The advent of the “pre-train, prompt” paradigm has recently extended its generalization ability and data efficiency to graph representation learning, following its achievements in Natural Language Processing (NLP). Initial graph prompt tuning approaches tailored specialized prompting functions for Graph Neural Network (GNN) models pre-trained with specific strategies, such as edge prediction, thus limiting their applicability. In contrast, another pioneering line of research has explored universal prompting via adding prompts to the input graph’s feature space, thereby removing the reliance on specific pre-training strategies. However, the necessity to add feature prompts to all nodes remains an open question. Motivated by findings from prompt tuning research in the NLP domain, which suggest that highly capable pre-trained models need less conditioning signal to achieve desired behaviors, we advocate for strategically incorporating necessary and lightweight feature prompts to certain graph nodes to enhance downstream task performance. This introduces a combinatorial optimization problem, requiring a policy to decide 1) which nodes to prompt and 2) what specific feature prompts to attach. We then address the problem by framing the prompt incorporation process as a sequential decision-making problem and propose our method, RELIEF, which employs Reinforcement Learning (RL) to optimize it. At each step, the RL agent selects a node (discrete action) and determines the prompt content (continuous action), aiming to maximize cumulative performance gain. Extensive experiments on graph and node-level tasks with various pre-training strategies in few-shot scenarios demonstrate that our RELIEF outperforms fine-tuning and other prompt-based approaches in classification performance and data efficiency. The code is available at https://github.com/JasonZhujp/RELIEF.
Graph neural networks, prompt tuning, reinforcement learning, hybrid action space, few-shot learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data miningccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Reinforcement learning
1. Introduction
Figure 1. A Comparison of Tuning Methods. Fine-tuning (upper left) updates the parameters of the pre-trained GNN model. Pre-training-dependent prompt tuning (right) freezes the GNN model and requires designing specialized prompt templates aligned with pre-training strategies, whereas feature prompt tuning (lower left) is applicable to any pre-training strategy.
[…]
Figure 1 Fully described in the text.
In recent years, Graph Neural Networks (GNNs) have been applied in diverse domains, including knowledge graphs (Liang et al., 2023), social networks (Cao et al., 2021), and recommender systems (Jiang et al., 2023), due to their powerful expressivity in capturing complex relationships in real-world data. However, traditional supervised learning on graphs struggles with limited labeled data and poor generalization to out-of-distribution samples (Yehudai et al., 2021). To overcome these challenges, significant efforts have focused on pre-trained GNN models (Wu et al., 2021; Xie et al., 2022), which are first trained on label-free graphs in a self-supervised manner to learn intrinsic properties, followed by fine-tuning on downstream tasks.
Despite advancements, the “pre-train, fine-tune” paradigm still exhibits drawbacks. The misalignment between pre-training and downstream objectives can lead to sub-optimal performance, known as negative transfer (Han et al., 2021). Moreover, under few-shot settings, pre-trained models are prone to overfitting, leading to catastrophic forgetting (Liu et al., 2021), which compromises their generalization ability. To tackle the issues, prompt learning, which has shown remarkable success in Natural Language Processing (NLP) (Brown et al., 2020; Liu et al., 2023b; Lester et al., 2021) and Computer Vision (CV) (Jia et al., 2022; Kirillov et al., 2023; Yao et al., 2023), emerges as a promising approach. By designing informative prompts to manipulate input data, prompt tuning aligns objectives between pretext and downstream tasks, thereby augmenting the potential of pre-trained models.
Naturally, researchers have extended the “pre-train, prompt” framework to graph domain (Sun et al., 2023b). Existing approaches are broadly divided into two categories based on whether they depend on pre-training strategies. For methods reliant on pre-training strategies, GPPT (Sun et al., 2022) and GraphPrompt (Liu et al., 2023a) unify pretext and downstream tasks as edge prediction. All in One (Sun et al., 2023a) introduces a prompt graph to match the graph-level contrastive learning used in pre-training. SGL-PT (Zhu et al., 2023) aligns the downstream task with masked node prediction using contrastive and generative learning objectives. However, these methods may fail if a GNN model is pre-trained with multiple self-supervised techniques through multi-task learning to enhance model’s capabilities, rather than solely simple link prediction.
In contrast, prompting methods agnostic to pre-training strategies offer greater compatibility. GPF (Fang et al., 2024), a pioneering method in this line, introduces graph feature-based prompt tuning, demonstrating that adding a learnable uniform feature prompt vector to each node is theoretically equivalent to any graph manipulation, without restricting pre-training strategies. GPF-plus (Fang et al., 2024), a variant of GPF, provides more flexible adaptation by inserting node-specific prompted features. SUPT (Lee et al., 2024), a subsequent study, innovated by assigning feature prompts at the subgraph level, captures intricate contextual nuances of each subgraph, leading to a decent performance improvement. Thus, feature-based prompting forms the basis of our paper due to its generality and effectiveness.
Nonetheless, it is critical to delve deeper into existing feature-based prompting methods. In NLP, soft prompt tuning adapts frozen language models for specific tasks using continuous prompt tokens (Liu et al., 2023b). Research shows that while increasing prompt token length involves more tunable parameters and thus enhances expressivity, excessively long prefixes can degrade test performance (Li and Liang, 2021; Zhu et al., 2022). In fact, T5-XXL can achieve strong results even with a single-token prompt, indicating that more capable models need less conditioning signal to exhibit desired behavior (Lester et al., 2021). This finding is even more enlightening in the context of graphs. Since most GNNs follow the message-passing mechanism, information from a feature prompt attached to a target node will propagate to its neighbors, diffusing the influence of prompting signals to a broader range.
We propose that given an adequately pre-trained GNN model, judiciously inserting minimal prompts as conditioning signals may suffice for downstream tasks. In the realm of feature prompting, “minimal” can be translated as an appropriate amount and magnitude of prompts. Conversely, incorporating prompts to an excessive number of nodes increases learnable parameters, raising the risk of overfitting (Li and Liang, 2021), particularly with few-shot downstream data. Moreover, since prompts are directly added to node features, large-magnitude prompts might overwhelm the original input’s feature space, causing pre-trained GNNs to perceive prompted graphs as unfamiliar, leading to a decline in transfer performance.
Hence, the key lies in strategically incorporating necessary and lightweight prompts into the original graph, allowing a GNN model to fully leverage its learned knowledge during pre-training without being overly disrupted, and to better generalize to downstream tasks through flexible feature prompts.
The choice of prompting on certain nodes and attaching specific feature prompts leads to a combinatorial optimization problem. To tackle this, we turn to Reinforcement Learning (RL), which effectively searches for heuristic strategies by training an agent (Mazyavkina et al., 2021), and propose a REinforcement LearnIng Empowered graph Feature prompting method, named RELIEF. We formulate the process of inserting prompts as a sequential decision-making problem. At each step, the RL agent: 1) selects which node to prompt (discrete action), and 2) determines the prompt content, i.e., specific values of the prompt vector (continuous action). Therefore, an RL algorithm capable of handling the discrete-continuous hybrid action space (Fan et al., 2019; Masson et al., 2016; Xiong et al., 2018) is employed. The prompted graph is then evaluated by the pre-trained GNN model. At the subsequent step, a new prompt generated by the RL agent is added to the previous prompted graph. Prompt addition and performance evaluation are repeated until a maximum step is reached. Our goal is to maximize the expected cumulative performance gain on the downstream task. Techniques like policy generalization (Ghosh et al., 2021; Wang et al., 2020; Raileanu and Fergus, 2021) are integrated to ensure training stability and effectiveness. Furthermore, we have designed two metrics: prompt coverage ratio and average prompt magnitude to quantify the extent to which the prompt affects the original input.
Our contributions are summarized as follows:
  • We suggest enhancing the performance of pre-trained GNN models on downstream tasks by adding necessary and lightweight feature prompts and design metrics to quantify their impact on inputs, offering a new perspective on graph prompt tuning.
  • As far as we know, we are the first to formulate feature prompt incorporation as a sequential decision-making problem, and propose RELIEF, an RL empowered graph feature prompting method optimized through trial and error in a hybrid action space.
  • Extensive experiments across both graph and node-level tasks with various pre-training strategies in few-shot scenarios demonstrate that RELIEF achieves superior classification performance and data efficiency compared to fine-tuning and other competitive prompt-based methods.
2. Preliminaries
Let an undirected graph instance be $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X})\in\mathbb{G}$, where $\mathcal{V}=\{v_{1},v_{2},\dots,v_{n}\}$ denotes the node set containing $n$ nodes; =$\mathcal{E}=\{(v_{i},v_{j})\mid v_{i},v_{j}\in\mathcal{V}\}$ denotes the edge set; $\mathbf{X}=\{x_{1},x_{2},\dots,x_{n}\}\in\mathbb{R}^{n\times D}$ denotes the node feature matrix, where $x_{i}\in\mathbb{R}^{1\times D}$ is the feature vector of node $v_{i}$ with feature dimension $D$. The adjacency matrix $\mathbf{A}\in\{0,1\}^{n\times n}$ is defined such that $\mathbf{A}_{ij}=1$ if $(v_{i},v_{j})\in\mathcal{E}$.
Fine-tuning and Graph Prompt Tuning.
Given a pre-trained GNN model $f_{\theta}$ and a task-specific projection head $g_{\phi}$ for making predictions, parameters $\theta$ and $\phi$ are fine-tuned on a downstream training dataset $\mathcal{D}=\{(\mathcal{G}_{1},y_{1}),\dots,(\mathcal{G}_{m},y_{m})\}$ to maximize the likelihood of correctly predicting label $y$ of graph $\mathcal{G}$, described as:
$\max_{\theta,\phi}\prod_{i=0}^{m}P_{f_{\theta},\,g_{\phi}}(y_{i}\mid\mathcal{G}_{i})=\max_{\theta,\phi}\prod_{i=0}^{m}g_{\phi}(f_{\theta}(\mathbf{X},\mathbf{A}))\left[y_{i}\right]$
where $\left[y_{i}\right]$ fetches the prediction probability of label $y_{i}$ from $g_{\phi}(\cdot)$. In contrast, the pre-trained GNN model $f_{\theta}$ remains frozen during prompt tuning. The primary focus shifts to learning a prompting function $u_{\psi}:\mathbb{G}\to\mathbb{G}$, which transforms the original graphs $\mathcal{G}$ to prompted graphs $u_{\psi}(\mathcal{G})$, used as inputs to the GNN model. The projection head $g_{\phi}$ is also tuned to coordinate with the prompting function. The optimal $u_{\psi}$ and $g_{\phi}$ are obtained by:
$\max_{\psi,\phi}\prod_{i=0}^{m}P_{f_{\theta},\,g_{\phi}}(y_{i}\mid u_{\psi}(\mathcal{G}_{i}))=\max_{\psi,\phi}\prod_{i=0}^{m}g_{\phi}(f_{\theta}(\mathbf{X}_{\psi},\mathbf{A}_{\psi}))\left[y_{i}\right]$
where $\mathbf{X}_{\psi}$ and $\mathbf{A}_{\psi}$ are prompted feature and adjacency matrix, respectively, acquired by prompting function $u_{\psi}$. For evaluation, the downstream testing graphs are first transformed by the prompting function and then passed through the pre-trained GNN model, followed by the projection head to produce predictions.
Graph Feature Prompt Tuning.
This line of prompting method introduces learnable components to the feature space of the input graph. Given the node feature matrix $\mathbf{X}=\{x_{1},x_{2},\dots,x_{n}\}$ of a graph $\mathcal{G}$, learnable prompt feature vectors $p_{1},p_{2},\dots,p_{n}\in\mathbb{R}^{1\times D}$ are added to node features. This can be viewed as the prompting function $u_{\psi}$, achieving a prompted feature matrix $\mathbf{X}^{\ast}$ expressed as:
$\mathbf{X}^{\ast}=\{x_{1}+p_{1},x_{2}+p_{2},\dots,x_{n}+p_{n}\}=\{x^{\ast}_{1},x^{\ast}_{2},\dots,x^{\ast}_{n}\}$
and $\mathbf{X}^{\ast}$ is then processed by the GNN model. Note that feature prompt tuning does not explicitly prompt on the adjacency matrix, thus the representation of a prompted graph is given by $f_{\theta}(\mathcal{G}^{\ast})=f_{\theta}(\mathbf{X}^{\ast},\mathbf{A})$. Different feature prompting methods vary in their prompt vector designs or learning approaches. A comparison of aforementioned tuning methods is depicted in Figure 1.
RL with Hybrid Action Space.
Most RL algorithms operate within either discrete or continuous action spaces  (Sutton and Barto, 2018). In our scenario, however, we deal with a hybrid action space  (Masson et al., 2016; Xiong et al., 2018), involving selecting nodes (discrete) and specifying feature prompts (continuous). Formally, discrete actions are chosen from a finite set $\mathcal{A}_{d}=\{a_{1},a_{2},\dots,a_{n}\}$, each corresponding to a real-valued continuous action $z\in\mathbb{R}^{D}$, where $D$ is the dimension of the continuous action space $\mathcal{A}_{c}=\{z_{1},z_{2},\dots,z_{n}\}$. H-PPO  (Fan et al., 2019) is an RL algorithm tailored for hybrid action spaces based on actor-critic architecture. It employs two parallel actor networks for discrete and continuous actions, respectively, alongside a single critic network for value estimation. H-PPO extends Proximal Policy Optimization (PPO)  (Schulman et al., 2017) by separately optimizing the two actors with an entropy-regularized PPO surrogate objective $\mathcal{L}^{\text{PPO}}$ to maximize expected cumulative discounted reward, while the critic minimizes Mean Squared Error (MSE) loss $\mathcal{L}^{\text{Critic}}$ to improve value estimation accuracy. For a detailed introduction to PPO and H-PPO, please refer to Appendix  E.
3. Method
In this section, we first introduce the formulation of a Markov Decision Process (MDP) designed for our feature prompting scenario (Section 3.1). We then detail the architecture of the policy network to instantiate the MDP (Section 3.2). Next, we present the overall framework of RELIEF, featuring alternating training (Section 3.3). To mitigate policy overfitting, we integrate policy generalization techniques (Section 3.4). Additionally, we describe two metrics designed to quantify the impact of the inserted feature prompts on original graphs (Section 3.5). Note that we use graph-level classification as an example throughout this section and will later extend to node tasks in the experiments section.
3.1. Incorporating Feature Prompts as MDP
In RL domain, the environment is typically modeled by an MDP (Sutton and Barto, 2018). With the aim of framing the process of incorporating prompts as an MDP, we present the design of the basic elements as follows.
Action Space.
Given a graph $\mathcal{G}$ with $n$ nodes, a discrete action $a$ is to select a node $v_{a}$ from the node set $\{v_{1},\dots,v_{n}\}$, and a continuous action $z\in\mathbb{R}^{1\times D}$ is to determine a real-valued vector for node $v_{a}$, where $D$ is the dimension of the initial node feature. Thus, a hybrid action, i.e., feature prompt, incorporated at time step $t$ can be denoted as $(a_{t},z_{t})=p^{a,z}_{t}$. We further define the prompt matrix at step $t$ as $\mathbf{P}_{t}=\{p_{1,t},\dots,p_{n,t}\}\in\mathbb{R}^{n\times D}$. $\mathbf{P}_{0}$ is initialized as a zero matrix, indicating no prompt is applied. Thus, the prompted feature $\mathbf{X}^{\ast}$ of a prompted graph $\mathcal{G}^{\ast}$ is obtained by $\mathbf{X}^{\ast}=\mathbf{X}+\mathbf{P}$.
State Transition.
We define the state space as node representations of the prompted graph $\mathcal{G}^{\ast}$ obtained from the pre-trained GNN model $f_{\theta}$. This allows the agent to be aware of node representations that closely tied to the GNN model, enabling precise control. Formally, the state at time step $t$ is denoted as:
$\begin{split}s_{t}&\coloneqq f_{\theta}(\mathcal{G}^{\ast}_{t-1})=f_{\theta}(\mathbf{X}^{\ast}_{t-1},\mathbf{A})=f_{\theta}(\mathbf{X}+\mathbf{P}_{t-1},\mathbf{A})\\ &=\{h^{\ast}_{1,t-1}\,,\dots,h^{\ast}_{n,t-1}\}\in\mathbb{R}^{n\times d}\end{split}$
where $h^{\ast}_{i,t-1}$ is the representation of node $v_{i}$ in the prompted graph $\mathcal{G}^{\ast}_{t-1}$, and $d$ is the dimension of the latent space. Notably, the current state builds upon the representation of the previous step. However, as different graphs contain varying numbers of nodes, this results in state spaces of different sizes, posing challenges for batch training. To fix the size, we set a maximum number of nodes $N$ and use zero padding, i.e., append $(N-n)$ zero vectors $\mathbf{0}\in\mathbb{R}^{1\times d}$ to the end of $h^{\ast}_{n,t}$, ensuring a fixed-size state representation:
(1) $\begin{split}s_{t}&\coloneqq f_{\theta}(\mathbf{X}^{\ast}_{t-1},\mathbf{A})\left|\right|\mathbf{0}_{(N-n)\times d}\\ &=\{h^{\ast}_{1,t-1},\dots,h^{\ast}_{n,t-1},\mathbf{0}_{n+1},\dots,\mathbf{0}_{N}\}\in\mathbb{R}^{N\times d}\end{split}$
where ||||| | denotes concatenation. When the agent executes action $p^{a,z}_{t}$ at step $t$, which inherently operates on node $v_{a}$’s feature prompt $p_{a}$, the prompted feature matrix then transits to $\mathbf{X}^{\ast}_{t}=\mathbf{X}+\mathbf{P}_{t}$, with $\mathbf{P}_{t}$ computed as:
$\mathbf{P}_{t}=\mathbf{P}_{t-1}+p^{a,z}_{t}=\{p_{1,t-1}\,,\dots,\,p_{a,t-1}+p^{a,z}_{t},\dots,\,p_{n,t-1}\}$
Finally, the next state $s_{t+1}$ is obtained by replacing $\mathbf{X}^{\ast}_{t-1}$ in Eq. (1) with $\mathbf{X}^{\ast}_{t}$. Practically, this involves feeding the prompted graph $\mathcal{G}^{\ast}_{t}$ into the pre-trained GNN model to derive new node representations, along with zero padding, to construct the subsequent state.
Reward Function.
An ideal reward function is goal-oriented, providing guiding signals about action values during exploration. While downstream task metrics such as AUC or F1-score might seem straightforward for graph classification tasks, we require a reward to measure the quality of prompt inserted at each step for each graph. However, these metrics are infeasible to apply to a single graph. Instead, we use the loss decrease as our instant reward, since it captures the concept of performance gain somewhat and can be derived from each graph. Formally, given two adjacent steps, the reward $r(s_{t},a_{t},z_{t},s_{t+1})$, or $r_{t}$ for short, is defined as:
(2) $r_{t}=\mathcal{L}_{t-1}-\mathcal{L}_{t}=\mathcal{L}\left(g_{\phi}\left(f_{\theta}(\mathcal{G}^{\ast}_{t-1})\right),y\right)-\mathcal{L}\left(g_{\phi}\left(f_{\theta}(\mathcal{G}^{\ast}_{t})\right),y\right)$
where $\mathcal{L}(\cdot)$ denotes a loss function associated with the downstream task, such as cross-entropy loss for classification. Hence, the reward is positive if the prompt assigned at this step leads to a loss decrease, and negative if it increases loss. Consequently, the cumulative reward reflects the total loss decrease from time step $1$ to $T$, serving as a surrogate measurement of overall performance improvement.
3.2. Policy Network Architecture
Figure 2. RELIEF pipeline. The policy network (upper) and the projection head (lower) are trained alternately. Feature prompts are incorporated during the agent sampling and prompt addition processes via discrete and continuous actors.
[…]
Figure 2 Fully described in the text.
RELIEF employs H-PPO, comprising two parallel actor networks and a single critic network collectively referred to as the policy network $\Pi_{\omega}$, where $\omega$ denotes the parameters of these networks.
In practice, all the three networks share the first few layers to encode state information. Since the state space is designed as the node representations of a prompted graph, we utilize the pre-trained GNN model $f_{\theta}$ as the state encoder. Thereafter, Multi-Layer Perceptrons (MLPs) with different output dimensions are connected to each of the three networks, facilitating the corresponding functions. Given a prompted graph $\mathcal{G}^{\ast}$, the forward propagation of the networks is presented as follows:
$\begin{split}p(a\mid s)&\leftarrow\textsc{Softmax}\left(\text{MLP}_{a}\left(f_{\theta}(\mathcal{G}^{\ast})\right)\right)\\ \boldsymbol{\mu}(s,a)&\leftarrow\text{MLP}_{z}\left(f_{\theta}(\mathcal{G}^{\ast})\right)\left[a\right]\\ V(s)&\leftarrow\text{MLP}_{c}\left(\textsc{Flatten}\left(f_{\theta}(\mathcal{G}^{\ast})\right)\right)\end{split}$
Let’s elaborate on each expression in turn.
Discrete Actor
Represents the discrete policy $\pi_{d}(a|s)$. Given node representations of a prompted graph as state $s$, i.e., $f_{\theta}(\mathcal{G}^{\ast})$ with size $N\times d$, the $\text{MLP}_{a}$ followed by a Softmax operation transforms $s$ into a discrete action probability $p(a|s)\in\mathbb{R}^{n}$. The agent then either samples a node $v_{a}$ based on this probability or greedily selects the node with the highest probability as a discrete action, indicating a stochastic or deterministic policy, respectively. Note that we set the entries in $p(a|s)$ associated with zero padding to zero, eliminating their chance of being selected, which reduces the number of valid discrete actions from $N$ to $n$.
Continuous Actor
Represents the continuous policy $\pi_{c}(z|s,a)$. Given state $s\in\mathbb{R}^{N\times d}$, the $\text{MLP}_{z}$ outputs $N$ parameters $\boldsymbol{\mu}\in\mathbb{R}^{1\times D}$ for all the $N$ possible discrete actions. Then $\boldsymbol{\mu}$ with index $\left[a\right]$ is chosen paired with the selected discrete action $a$. Subsequently, the agent constructs a Gaussian distribution based on $(\boldsymbol{\mu},\boldsymbol{\sigma})$ and samples a vector $z\in\mathbb{R}^{1\times D}$ as a prompt feature $p^{a,z}$ stochastically, or directly use $\boldsymbol{\mu}$ as the action deterministically, where the standard deviation $\boldsymbol{\sigma}\in\mathbb{R}^{1\times D}$ can either be learned or pre-defined. To ensure $p^{a,z}$ remains within a desired range, each dimension of $z$ is clamped to range $\left[-z_{\text{max}},z_{\text{max}}\right]$, where $z_{\text{max}}>0$ is a hyper-parameter controlling the scale of the prompt added at every step.
Critic
Used for estimating the state-value function. At its core, it maps a state $s$ to a real value $V(s)\in\mathbb{R}$. However, there exists a dimension discrepancy: our state space possesses a node-level granularity, whereas value estimation is based on a global view. Therefore, we use Flatten operation to transform the size of state $s$ from $N\times d$ to $1\times Nd$. The flattened vector is then passed through $\text{MLP}_{c}$ to produce a real value, treated as the estimation of $V(s)$.
It is worth noting that the state encoder in the policy network $\Pi_{\omega}$ is a direct copy of the pre-trained GNN model and keeps frozen during policy learning. This implies it is by updating the parameters of the MLPs, that the actors are enabled to map states to actions, and the critic is enabled to map states to state values. Such structure has been widely used in RL from Human Feedback (RLHF) practice (Ouyang et al., 2022; Dai et al., 2023), where reward models and policy networks are constructed on fixed LLMs with learnable linear layers appended.
3.3. Overall Framework of RELIEF
RELIEF contains two trainable modules: the policy network and the projection head, as illustrated in Figure 2. Effective coordination of these two modules greatly enhances the prompting performance. We train them alternatively in practice, as described below.
3.3.1. Policy network training
Given the frozen pre-trained GNN model $f_{\theta}$, the policy network $\Pi_{\omega}$, the projection head $g_{\phi}$ and a graph $\mathcal{G}$ with label $y$ containing $n$ nodes, the initial loss $\mathcal{L}_{0}$ is computed via $\mathcal{L}(g_{\phi}(f_{\theta}(\mathcal{G})),y)$, with no prompt attached.
At each step, the agent samples a feature prompt $p^{a,z}_{t}$ with respect to the policies $\pi_{d}$ and $\pi_{c}$, and adds it to node $v_{a}$. The prompted graph is then passed to the GNN model followed by the projection head to obtain the current loss according to label $y$. By comparing the current loss with the previous one, an instant reward is computed by Eq. (2). As a result, the agent collects a transition, denoted as a tuple $(s,a,z,r,s^{\prime})$. The prompt addition process is repeated $n$ times, equivalent to the number of nodes, theoretically giving every node a chance of being prompted. Notably, both actors use stochastic policies for better exploration during training.
Next, the collected $n$ transitions are used to update the policy network. The two actors are independently trained using the PPO surrogate objective $\mathcal{L}^{\text{PPO}}$, whereas the critic is trained by MSE loss $\mathcal{L}^{\text{Critic}}$. The above process is batchified by consuming a batch of graphs simultaneously to increase sampling and training efficiency.
3.3.2. Projection head training
This aligns predictions with their correct labels by coordinating the projection head with representations of the ultimate prompted graphs. Thus, we use the learned policy to insert feature prompts for $n$ steps to obtain the final prompted graph $\mathcal{G}^{\ast}$. Here, we modify the continuous policy to be deterministic, ensuring identical prompt vector values for the same state and discrete action. This guarantees stable prompted graphs and thus consistent representations, which together with the labels, are used to supervise the projection head update. Given $m$ graph samples, the projection head $g_{\phi}$ is updated by minimizing the objective:
(3) $\min_{\phi}{\frac{1}{m}\sum^{m}_{i=1}\mathcal{L}\left(g_{\phi}\left(f_{\theta}(\mathcal{G}^{\ast}_{i})\right),y\right)}$
where $\mathcal{L}(\cdot)$ is the same loss function as in Eq. (2). To facilitate a more rapid coordination of the projection head with the policy, the projection head is updated $q$ times.
To summarize, the above alternating process – training policy network once and projection head $q$ times – defines a training epoch. At the evaluation stage, we directly apply the two learned actors to incorporate feature prompts step by step into the downstream validation and testing graphs. These prompted graphs are transformed by the pre-trained GNN model and the learned projection head to produce predictions, which are then measured by downstream metrics.
3.4. Policy Generalization
When trained in a limited set of environments, general RL algorithms often exhibit overfitting, resulting in poor generalization to unseen testing scenarios (Wang et al., 2020; Raileanu and Fergus, 2021), which is exactly the case in few-shot prompt tuning. To tackle this, we introduce a policy generalization technique, named LEEP (Ghosh et al., 2021), which can be seamlessly integrated with PPO, and thus compatible to our RELIEF.
Practically, LEEP is an ensemble-based method designed for discrete action spaces, adding a regularization term to PPO surrogate objective $\mathcal{L}^{\text{PPO}}$ for updating the actor network. LEEP learns a generalized joint policy by leveraging all sub-policies. To generalize our discrete policy $\pi_{d}(a|s)$, we learn $l$ discrete sub-policies $\{\pi_{d,1},\dots,\pi_{d,l}\}$. Each $\pi_{d,i}$ gathers transitions from a training graph set $\mathcal{D}_{i}$, which is bootstrap-sampled from the entire training set $\mathcal{D}$. Each $\pi_{d,i}$ is updated by maximizing the expected reward, while minimizing the disagreement between $\pi_{d,i}$ itself and the joint discrete policy $\pi_{d,\,J}$, i.e., maximizing the following objective:
(4) $\mathcal{L}_{d,i}=\mathcal{L}^{\text{PPO}}_{d,i}-\alpha_{d}\mathbb{E}_{s\sim\pi_{d,i},\,\mathcal{D}_{i}}\left[D_{\text{KL}}\Big{(}\pi_{d,i}(a|s)\left|\right|\pi_{d,\,J}(a|s)\Big{)}\right]$
where $\alpha_{d}>0$ is a uniform penalty hyper-parameter for all discrete sub-polices and $D_{\text{KL}}$ denotes the KL-divergence. The discrete joint policy $\pi_{d,\,J}$ combines all sub-policies $\pi_{d,i}$, and is computed as:
(5) $\pi_{d,\,J}(a|s)=\frac{\max_{i=1,\dots,l}\pi_{d,i}(a|s)}{\sum_{a^{\prime}}\max_{i=1,\dots,l}\pi_{d,i}(a^{\prime}|s)}$
indicating that to obtain the discrete action probability $p(a|s)$ given by $\pi_{d,\,J}$, we take the maximum probability across all $l$ sub-policies for each action $a$, and normalize these maxima by their sum across all actions.
Since RELIEF requires hybrid action spaces, we extend LEEP to continuous action spaces. Similarly, we learn $l$ continuous sub-policies $\{\pi_{c,1},\dots,\pi_{c,l}\}$, each exploring in its respective training contexts $\mathcal{D}_{i}$. In other words, we employ $l$ parallel H-PPO algorithms, but still with a single critic. Each continuous sub-policy $\pi_{c,i}$ is trained by maximizing the following objective:
(6) $\mathcal{L}_{c,i}=\mathcal{L}^{\text{PPO}}_{c,i}-\alpha_{c}\mathbb{E}_{s\sim\pi_{c,i},\,\mathcal{D}_{i}}\left[D_{\text{KL}}\Big{(}\pi_{c,i}(a|s)\left|\right|\pi_{c,\,J}(a|s)\Big{)}\right]$
where $\alpha_{c}>0$ is a uniform penalty hyper-parameter for continuous sub-policies. The continuous joint policy $\pi_{c,\,J}$ is defined as:
(7) $\pi_{c,\,J}(z|s,a)=\frac{1}{l}\sum_{i=1}^{l}\pi_{c,i}(z|s,a)=\frac{1}{l}\sum_{i=1}^{l}\boldsymbol{\mu}_{i}(s,a)$
indicating that the average mean $\boldsymbol{\mu}_{i}$ of all sub-policies is used as the continuous joint policy.
To sum up, the policy network, enhanced by policy generalization techniques, now comprises $l$ discrete actors, $l$ continuous actors and a single critic. During training, discrete and continuous actors within each pair are independently updated using Eq. (4) and (6), respectively, followed by the critic update using $\mathcal{L}^{\text{Critic}}$. Each pair of actors is sequentially updated. The joint policies $\pi_{d,\,J}$ and $\pi_{c,\,J}$ are applied to incorporate prompt features when training the projection head or testing at the evaluation stage. The pseudo-code for RELIEF can be found in Appendix F. Additionally, a detailed complexity analysis is in Appendix G.
Table 1. ROC-AUC (%) and standard deviation for graph classification on molecule property prediction benchmark under 50-shot scenario with various pre-training and tuning strategies. The best results for each dataset and pre-training strategy are highlighted in bold, and the runner-up is underlined.
Tuning Strategy BBBP Tox21 ToxCast SIDER ClinTox MUV HIV BACE Avg.
Infomax FT 65.26 $\pm$1.05 71.54 $\pm$0.73 57.98 $\pm$0.42 53.45 $\pm$0.49 59.01 $\pm$2.18 64.08 $\pm$0.84 65.13 $\pm$2.50 64.57 $\pm$0.71 62.63
GPF 66.30 $\pm$0.94 71.26 $\pm$0.30 58.33 $\pm$0.27 53.65 $\pm$0.34 63.11 $\pm$2.35 67.19 $\pm$0.10 64.09 $\pm$0.66 64.32 $\pm$0.88 63.53
GPF-plus 66.12 $\pm$1.27 71.54 $\pm$0.63 58.43 $\pm$0.27 53.76 $\pm$1.15 65.06 $\pm$1.56 67.20 $\pm$0.17 64.81 $\pm$1.62 65.88 $\pm$1.72 64.10
SUPTsoft 66.45 $\pm$0.85 71.82 $\pm$0.16 58.71 $\pm$0.44 53.72 $\pm$0.45 65.96 $\pm$3.37 66.52 $\pm$0.55 65.51 $\pm$2.31 66.26 $\pm$2.45 64.37
SUPThard 66.14 $\pm$0.85 71.55 $\pm$0.37 58.65 $\pm$0.33 53.82 $\pm$0.49 65.90 $\pm$3.41 66.53 $\pm$0.35 65.59 $\pm$2.26 66.40 $\pm$0.70 64.32
RELIEF 67.93 $\pm$0.73 71.58 $\pm$0.13 58.78 $\pm$0.17 53.95 $\pm$0.40 66.28 $\pm$2.07 67.23 $\pm$0.08 67.00 $\pm$0.55 67.95 $\pm$0.52 65.09
AttrMasking FT 66.48 $\pm$0.44 72.32 $\pm$0.19 57.35 $\pm$0.42 54.62 $\pm$0.58 59.71 $\pm$4.43 61.49 $\pm$1.04 58.37 $\pm$1.51 65.41 $\pm$2.09 61.97
GPF 66.67 $\pm$0.60 72.31 $\pm$0.30 58.01 $\pm$0.27 55.65 $\pm$1.92 67.25 $\pm$2.41 62.58 $\pm$0.31 58.71 $\pm$2.56 68.54 $\pm$1.24 63.72
GPF-plus 66.29 $\pm$0.36 72.75 $\pm$0.38 57.91 $\pm$0.38 55.05 $\pm$1.20 68.89 $\pm$2.86 62.70 $\pm$0.48 58.71 $\pm$1.07 68.07 $\pm$1.20 63.80
SUPTsoft 66.28 $\pm$0.81 72.76 $\pm$0.41 58.28 $\pm$0.34 54.70 $\pm$1.03 68.58 $\pm$4.54 62.81 $\pm$0.25 58.95 $\pm$1.71 66.41 $\pm$0.49 63.60
SUPThard 66.93 $\pm$0.91 72.75 $\pm$0.41 58.18 $\pm$0.44 55.20 $\pm$1.26 68.36 $\pm$3.57 62.81 $\pm$0.22 59.02 $\pm$1.61 68.92 $\pm$0.96 64.02
RELIEF 67.18 $\pm$0.55 72.40 $\pm$0.32 58.41 $\pm$0.14 56.54 $\pm$0.75 74.39 $\pm$1.18 63.18 $\pm$0.17 59.42 $\pm$0.55 69.37 $\pm$0.70 65.11
ContextPred FT 62.82 $\pm$0.83 70.11 $\pm$0.38 57.68 $\pm$0.69 56.68 $\pm$0.78 60.99 $\pm$2.28 61.28 $\pm$1.93 59.14 $\pm$1.93 64.47 $\pm$2.47 61.65
GPF 61.65 $\pm$0.76 70.42 $\pm$0.29 58.51 $\pm$0.38 56.55 $\pm$0.46 58.80 $\pm$2.57 64.83 $\pm$0.32 59.68 $\pm$2.20 68.57 $\pm$1.19 62.38
GPF-plus 61.25 $\pm$0.73 70.12 $\pm$0.42 57.64 $\pm$0.70 56.86 $\pm$0.52 61.43 $\pm$3.06 64.59 $\pm$0.28 59.15 $\pm$2.23 68.58 $\pm$1.42 62.45
SUPTsoft 61.85 $\pm$1.56 70.37 $\pm$0.14 57.82 $\pm$0.37 56.57 $\pm$0.50 59.13 $\pm$1.99 64.59 $\pm$1.00 56.58 $\pm$1.94 68.92 $\pm$0.92 61.98
SUPThard 62.29 $\pm$1.41 70.40 $\pm$0.15 58.06 $\pm$0.39 56.34 $\pm$0.72 59.18 $\pm$2.07 64.57 $\pm$0.69 59.08 $\pm$1.47 69.40 $\pm$1.62 62.42
RELIEF 62.99 $\pm$0.67 70.51 $\pm$0.14 58.58 $\pm$0.04 56.89 $\pm$0.25 63.06 $\pm$1.74 64.87 $\pm$0.31 59.72 $\pm$1.67 70.19 $\pm$0.96 63.35
GCL FT 62.13 $\pm$1.66 61.35 $\pm$0.88 53.96 $\pm$0.80 52.63 $\pm$0.71 68.73 $\pm$4.74 51.96 $\pm$1.93 58.67 $\pm$1.41 49.94 $\pm$3.80 57.42
GPF 61.58 $\pm$1.81 59.92 $\pm$1.29 54.44 $\pm$0.31 51.21 $\pm$0.56 75.52 $\pm$3.08 52.26 $\pm$1.28 58.37 $\pm$1.16 56.09 $\pm$2.44 58.67
GPF-plus 62.19 $\pm$1.45 60.13 $\pm$0.52 54.43 $\pm$0.43 50.90 $\pm$0.78 75.50 $\pm$1.11 52.40 $\pm$1.55 58.30 $\pm$1.68 59.55 $\pm$3.17 59.18
SUPTsoft 63.96 $\pm$0.85 60.13 $\pm$0.52 54.57 $\pm$0.76 51.44 $\pm$1.24 78.03 $\pm$2.32 51.86 $\pm$1.48 59.30 $\pm$2.59 57.39 $\pm$2.64 59.59
SUPThard 64.15 $\pm$0.96 60.56 $\pm$0.20 54.72 $\pm$0.71 51.60 $\pm$0.85 77.88 $\pm$2.06 51.78 $\pm$1.72 59.24 $\pm$2.42 59.68 $\pm$1.46 59.95
All in One 62.90 $\pm$3.77 61.49 $\pm$0.96 54.72 $\pm$1.03 52.73 $\pm$0.84 66.59 $\pm$3.84 53.34 $\pm$4.20 60.60 $\pm$3.80 59.57 $\pm$5.62 58.99
RELIEF 64.80 $\pm$0.69 61.45 $\pm$0.35 55.03 $\pm$0.39 52.82 $\pm$0.59 79.75 $\pm$1.28 53.52 $\pm$2.17 60.44 $\pm$2.39 62.97 $\pm$1.19 61.35
3.5. Metrics for Quantifying Prompts Impact
Our RELIEF stands out from existing feature prompting methods (Fang et al., 2024; Lee et al., 2024) by incorporating necessary and lightweight feature prompts. To measure the disturbance to the original input space caused by these prompts, we introduce two metrics: Prompt Coverage Ratio (PCR) and Average Prompt Magnitude (APM).
Prompt Coverage Ratio.
Given a graph $\mathcal{G}$ with $n$ nodes, and the final prompt matrix $\mathbf{P}=\{p_{1},\dots,p_{n}\}\in\mathbb{R}^{n\times D}$ obtained by adding feature prompts $p^{a,z}$ over $n$ steps, starting from a zero matrix, PCR is calculated as:
(8) $\text{PCR}(\mathcal{G})=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\left[p_{i}\neq\mathbf{0}_{1\times D}\right]\;\in\left[0,1\right]$
where $\mathbf{1}\left[\cdot\right]$ is the indicator function, which equals 1 if the feature prompt $p_{i}$ is not a zero vector $\mathbf{0}_{1\times D}$ (i.e., a valid prompt), otherwise 0. PCR signifies the proportion of nodes prompted at least once during the entire prompt addition process.
Average Prompt Magnitude.
We use L1-norm of the valid feature prompts averaged over dimensions, to describe the magnitude of inserted prompts, which is computed as:
(9) $\text{APM}(\mathcal{G})=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{D}\mathbf{1}\left[p_{i}\neq\mathbf{0}_{1\times D}\right]\cdot\left\|p_{i}\right\|_{1}\;\in\left[0,+\infty\right)$
where $\left\|p_{i}\right\|_{1}$ denotes the Manhattan norm of $p_{i}$, i.e., the sum of the absolute values of all entries of $p_{i}$. Thus, APM characterises the scale of valid feature prompts, offering a tangible measurement in terms of “lightweight”.
PCR and APM collectively assess how extensively and significantly feature prompts are integrated, which are applicable for evaluating any feature prompting methods.
4. Related Work
Figure 3. Tuning process of RELIEF on BACE. (a) and (b) present the ROC-AUC and average reward curves, respectively, for the training, validation, and testing sets across 5 random seeds. (c) and (d) illustrate the reward distributions for the training and testing sets as the tuning epochs progress, with white vertical lines indicating the average rewards for each distribution.
[…]
Figure 3 Fully described in the text.
4.1. Graph Prompt Tuning
Graph prompt tuning research can be categorized based on whether the prompting methods are independent of pre-training strategies. For methods paired with specific pre-trained GNN models, GPPT  (Sun et al., 2022) and GraphPrompt  (Liu et al., 2023a) align pretext and downstream tasks using link prediction. All in One  (Sun et al., 2023a) unifies tasks at graph level and incorporates learnable prompt graphs tuned by meta-learning. PSP (Ge et al., 2024) and Self-Pro (Gong et al., 2024) further adapt prompting approaches to heterophilous graphs by introducing graph-level dual-view and asymmetric contrastive learning, respectively, as pretext tasks, preserving more graph structural information.
Although these methods have shown success, their reliance on specific pre-training strategies limits their applicability. In contrast, feature prompt tuning, has demonstrated comparable results across various pre-training strategies. Pioneering works like GPF, GPF-plus  (Fang et al., 2024), and SUPT  (Lee et al., 2024) introduce learnable feature prompts to every node. However, they potentially overemphasize prompts and disrupt the input space, which may degrade performance.
To the best of our knowledge, our work is the first to apply RL to graph feature prompt tuning, aiming to optimize a policy for inserting necessary and lightweight feature prompts to enhance the downstream tasks performance of pre-trained GNN models.
4.2. RL for Graph Representation Learning
The integration of RL into graph representation learning has advanced significantly. MAG-GNN  (Kong et al., 2024) uses DQN to select optimal subgraph subsets to achieve high expressivity. WSD  (Wang et al., 2023) employs weighted sampling for subgraph counting, with edge weights determined by DDPG. SUGAR  (Sun et al., 2021) preserves hierarchical graph properties by selecting subgraphs using an RL pooling mechanism based on Q-learning. GPA  (Hu et al., 2020) utilizes DQN to learn optimal annotation strategies for valuable nodes in active search, while GraphCBAL  (Yu et al., 2024) further employs A2C for a class-balanced active search strategy.
While these works typically apply basic RL algorithms to either select actions from finite sets or determine single real values, RELIEF employs RL in a hybrid action space, skillfully integrating policy generalization techniques to achieve intricate control, representing a more cohesive fusion of RL and graph learning.
5. Experiments
5.1. Few-shot Graph Classification
GNN Architecture and Datasets.
We use a 5-layer GIN (Xu et al., 2018) as the base architecture for the GNN model, which is pre-trained on chemistry datasets (Hu et al., 2019) and prompt tuned on molecule properties prediction benchmark from MoleculeNet  (Wu et al., 2018) as downstream tasks, following the setup in existing feature prompting works (Fang et al., 2024; Lee et al., 2024). Detailed descriptions of these datasets are provided in Appendix B.1.
Pre-training Strategies.
To demonstrate the generality of our RELIEF, we pre-train the GIN model with four common strategies at both node and graph levels: Deep Graph Infomax (Infomax) (Veličković et al., 2018), Attribute Masking (AttrMasking) (Hu et al., 2019), Context Prediction (ContextPred) (Hu et al., 2019), and Graph Contrastive Learning (GCL) (You et al., 2020). Detailed descriptions of these methods are available in Appendix B.2.
Baselines.
We compare RELIEF with the following tuning approaches, all of which involve adjusting the projection head: additionally, Fine-Tuning (FT) tunes the parameters of the pre-trained GNN model; feature prompt tuning methods, including GPF, GPF-plus (Fang et al., 2024), SUPTsoft and SUPThard (Lee et al., 2024), tune the parameters of learnable feature prompts with the GNN model fixed; other prompt-based method, All in One (Sun et al., 2023a), uses a frozen GNN model specifically pre-trained by GCL, adapting learnable graph prompts, is included as it strictly aligns with our experimental setting.
Training Scheme.
Since molecular predictions are multi-label binary classifications, we use binary cross-entropy loss for the reward function and evaluate by averaging the ROC-AUC of each label. We construct few-shot scenarios with 50 samples for prompt tuning. We leverage early stopping based on validation set performance, unlike previous works that report results from the last epoch. Five rounds of experiments are conducted with different random seeds. The number of sub-policies $l$ is set to 3, with penalty factors $\alpha_{d}$ selected from {1e0, 1e1, 1e3, 1e5} and $\alpha_{c}$ set to 0.1. The maximum prompt scale per step $z_{\text{max}}$ is selected from $\{0.1,0.5,1.0\}$. The projection head is chosen from $\{1,2,3\}$-layer MLPs and trained $q$ times per epoch, ranging from $\{1,2,3\}$. Detailed data splitting and hyper-parameter settings are provided in Appendix B.3. All the experiments were conducted on an NVIDIA A800 PCIe 80GB GPU.
Table 2. PCR, APM and overall impact (OV.) of feature prompting methods averaged on all graph classification task.
Tuning Strategy PCR APM ($10^{-2}$) OV. ($10^{-2}$) ROC-AUC
GPF 1.00 7.15 7.15 62.08
GPF-plus 1.00 6.69 6.69 62.38
SUPTsoft 1.00 6.46 6.46 62.39
SUPThard 0.65 6.10 3.97 62.68
RELIEF 0.61 6.03 3.68 63.72
Performance Results.
Based on the results in Table 1, our RELIEF achieves superior graph classification performance in few-shot settings, surpassing baselines in 28/32 tasks, with an average improvement of 1.64% over fine-tuning and 1.04% over the runner-up. Notably, RELIEF is the only method consistently outperforming fine-tuning across all tasks. Moreover, we observe that in All in One, inserting prompt graphs with fewer nodes yields better results. This suggests that compact prompt structures may enhance prompt tuning, indirectly supporting our motivation.
We illustrate the tuning process of RELIEF on the BACE dataset pre-trained with Infomax in Figure 3. The ROC-AUC curve demonstrates a smooth and monotonically increasing trend, indicating the stability of the training process. The reward curve and distributions show reward improvement as the epochs progress, confirming the effectiveness of RELIEF’s agent in enhancing downstream task performance. Detailed tuning process analysis is in Appendix B.4.
Furthermore, we measure the impact of prompts on input graphs in terms of PCR and APM, and further multiply them to express the overall impact (OV.). Results are averaged across all tasks, as outlined in Table 2. RELIEF exhibits the smallest PCR, APM and overall impact compared to baselines. Interestingly, SUPThard also prompts on specific nodes with a PCR of 0.65, as each basis prompt vector affects only the top-ranked nodes, where the top-rank ratio is a hyper-parameter, making it a rule-based node selection method. In contrast, RELIEF does not rely on priori information, such as the usage of basis vectors or a top-rank ratio, thereby offers a highly flexible prompting policy based on the actors. Notably, the flexibility of graph manipulations is crucial for prompting performance (Sun et al., 2023a, b), contributing to RELIEF’s superior results.
5.2. Data Efficiency
Table 3. The minimum proportion (%) of data required for surpassing the performance of full-shot FT. The minimum proportion are highlighted in bold. $\boldsymbol{\times}$ denotes cannot surpass.
Methods BBBP Tox21 SIDER Cinltox Bace
Infomax GPF $\times$ 60 60 $\times$ 55
GPF-plus $\times$ 60 60 65 50
SUPTsoft $\times$ 65 60 70 45
SUPThard $\times$ 60 55 65 50
RELIEF 75 50 40 60 40
ContextPred GPF $\times$ 55 55 70 95
GPF-plus $\times$ 50 45 60 90
SUPTsoft $\times$ 65 50 65 85
SUPThard $\times$ 65 50 75 85
RELIEF 90 50 40 55 65
Figure 4. Performance of feature prompting methods with training data scaling up, compared to full-shot FT.
[…]
Figure 4 fully described in the text.
To evaluate the data efficiency of feature prompting methods, we incrementally add 5% of the training samples until the average ROC-AUC over five random seeds surpasses full-shot fine-tuning performance. We tune hyper-parameters after each data increment to optimize performance. We report the minimum proportion of data required for each method to exceed fine-tuning on five datasets pre-trained with Infomax and ContextPred strategies.
Table 3 shows that RELIEF requires the least data for all tasks, while other baselines, more or less, fail to surpass full-shot fine-tuning even with the entire dataset. Figure 4 illustrates the performance trends on BBBP dataset pre-trained with Infomax, which demonstrates that RELIEF’s overall performance presents the most noticeable improvement as data scales up, whereas other methods fluctuate and ultimately fail to exceed fine-tuning.
We attribute RELIEF’s superior data efficiency to its RL-based learning paradigm, which posses two advantages. Firstly, it reduces learning difficulty by optimizing one feature prompt at a time, allowing the agent to focus on a single state-to-action mapping and avoid complex node interactions that baseline methods face when optimizing entire feature prompts at once. Secondly, the step-by-step insertion and evaluation of prompts expose the agent to various state patterns and loss values, acting as data augmentation and providing additional supervision signals in few-shot settings.
5.3. Additional Experiments
Due to space constraints, we relocate additional experiments to the Appendices. Here, we provide an overview to highlight their critical conclusions.
Node Classification (Appendix  A). We extend RELIEF to node-level tasks by adding prompts to the $k$-hop subgraph induced by the target node (Liu et al., 2023a; Sun et al., 2023a). Experiments demonstrate RELIEF’s top-tier performance across multiple datasets compared to baselines. A case study is conducted to reveal that RELIEF can achieve its full potential when given adequately pre-trained GNN models.
Ablation Study (Appendix  C). We compare RELIEF against its three variants: random discrete policy, random continuous policy, and linear probing (solely training the projection head) (Kumar et al., 2022) on node and graph-level tasks. Results indicate significant performance drops when either policy is deactivated or feature prompting components are discarded.
Parameter Analysis (Appendix  D). We examine the sensitivity of key hyper-parameters of RELIEF including the maximum feature prompt scale per step $z_{\text{max}}$, number of sub-policies $l$, and penalty factors $\alpha_{d}$ and $\alpha_{c}$. We also validate the necessity of using policy generalization techniques. The findings confirm RELIEF’s robust performance across varied hyper-parameter settings.
5.4. Why RELIEF works?
Despite of RELIEF’s superior performance in graph and node-level tasks, it is essential to explain why RELIEF, utilizing an RL framework, fulfills our motivation – achieving powerful downstream performance by adding necessary and lightweight prompts.
Powerful. RL demonstrates significant efficacy in solving combinatorial optimization problems (Mazyavkina et al., 2021; Kong et al., 2024). With a goal-oriented reward function and adequate exploration within stable state and reward dynamics, the RELIEF agent identifies an optimal prompting policy to maximize performance gains.
Necessary. The agent makes decisions based on state patterns and learned policies, displaying specific node selection preferences when faced with similar prompted features. Suboptimal node selection at any step leads to a cumulative reward decrease, which is captured by the critic and rectified through policy updates, thereby avoiding unnecessary prompts.
Lightweight. As the case study analyzed in Appendix A.6, RELIEF tends to preserve pre-trained graph knowledge rather than overwhelm it. Larger-scale prompts cause the GNN model to generate unfamiliar representations, disrupting coordination between representations and the projection head, potentially leading to negative rewards. These punishments are applied to the actor, urging it to provide prompts that avoid excessive disturbance to the initial input through policy updates.
6. Conclusion
In this paper, inspired by the marginal effect of increasing prompt token length on performance improvement in LLMs, we propose RELIEF, an RL empowered graph feature prompt tuning method. RELIEF enhances the downstream performance of pre-trained GNN models by incorporating only necessary and lightweight feature prompts to original graphs. To tackle the combinatorial optimization problem, we employ an RL algorithm designed for hybrid action spaces and integrated policy generalization techniques. Extensive experiments on graph and node-level tasks demonstrate RELIEF’s effectiveness and data efficiency compared to fine-tuning and other prompt-based methods.
Acknowledgements.
This work is supported by National Natural Science Foundation of China No. 62202172.
Appendix A Few-Shot Node Classification
A.1. Extend to Node-Level Tasks
Given that feature prompting methods are originally designed to address graph-level tasks (Fang et al., 2024), adapting them to node-level tasks requires careful consideration.
Fortunately, prior research has successfully unified various levels of tasks as subgraph-level tasks (Liu et al., 2023a; Sun et al., 2023a). This unification is achieved by generating induced graphs for nodes or edges, which enables feature prompting methods to operate on $k$-hop subgraphs. These subgraphs act as proxies for the target nodes or edges during prompt tuning and downstream tasks evaluation. Here, we examine the capability of RELIEF on node classification tasks, alongside other existing feature prompting approaches.
A.2. Details of Datasets
We adopt Cora, Citeseer and Pubmed (Kipf and Welling, 2017) and Amazon-Co-Buy (Computer and Photo) (Shchur et al., 2018) for node classification. Cora, Citeseer, and Pubmed are citation networks, where each node represents a publication with a sparse bag-of-words feature vector, and edges signify citation links. Computer and Photo are segments of the Amazon co-purchase graph, with nodes representing products, edges indicating frequent co-purchases, node features encoded as bag-of-words from product reviews, and class labels derived from product categories. The statistics of these dataset are summarized in Table 4.
A.3. Details of Pre-training Strategies
To complement the node and graph-level pre-training strategies used in graph classification experiments, we focus on edge-level pre-training strategies to demonstrate the generality of RELIEF. Therefore, we include all previously mentioned feature prompting methods and additionally involve GPPT (Sun et al., 2022) and GraphPrompt (Liu et al., 2023a) as baselines, which are tailored to GNN models pre-trained at edge-level and capable of transferring to node classification.
Formally, GPPT uses masked edge prediction (MaskedEdge) as pretext task, which is optimized by a binary cross-entropy loss $\mathcal{L}(\cdot)$ formulated as:
$\min_{\theta}\sum_{v_{i},v_{j}}\mathcal{L}\left(f_{\theta}(v_{i})^{\top}f_{\theta}(v_{j}),\mathbf{A}_{ij}\right)$
where $\mathbf{A}_{ij}$ represents if node $vi$ and $vj$ are connected given graph $\mathcal{G}$ with the adjacency matrix $\mathbf{A}$. Moreover, GraphPrompt employs contrastive learning using positive and negative node pairs determined based on edge connectivity (ContraEdge). The pre-training loss is defined as follows:
$\begin{split}\min_{\theta}-\frac{1}{\left|\mathcal{V}\right|}&\sum_{v\in\mathcal{V}}\frac{1}{\left|\mathcal{V}^{+}\right|}\cdot\\ &\sum_{v^{+}\in\mathcal{N}(v)}\log\frac{\exp(h_{v}^{\top}h_{v^{+}}/\tau)}{\exp(h_{v}^{\top}h_{v^{+}}/\tau)+\sum_{v^{-}\in\mathcal{V}^{-}}\exp(h_{v}^{\top}h_{v^{-}}/\tau)}\end{split}$
where $h_{v}$, $h_{v^{+}}$ and $h_{v^{-}}$ are representations of node $v$, $v^{+}$ and $v^{-}$ produced by $f_{\theta}(\mathcal{G})$. $\mathcal{V}^{+}$ is the positive node set containing the neighbors of node $v$, while $\mathcal{V}^{-}$ is the negative node set constructed by random sampling. $\tau$ is the temperature hyper-parameter.
Table 4. Statistics of the datasets used in our node classification experiments.
Dataset # Nodes # Edges # Features # Classes
Cora 2708 5429 1433 7
CiteSeer 3327 9104 3703 6
PubMed 19717 88648 500 3
Computers 13752 491722 767 10
Photo 7650 238162 745 8
Table 5. Accuracy (%) and Macro F1-score (%) with respective standard deviation for node classification under 10-shot scenario with various two edge-level pre-training and various tuning strategies. The best results for each dataset and pre-training strategy are highlighted in bold, and the runner-up is underlined.
Tuning Strategy Cora CiteSeer PubMed Computers Photos
 Accuracy Macro F1 Accuracy Macro F1 Accuracy Macro F1 Accuracy Macro F1 Accuracy Macro F1
MaskedEdge FT  54.40 $\pm$1.44 53.27 $\pm$1.45 61.28 $\pm$0.93 56.60 $\pm$0.92 71.40 $\pm$1.61 67.11 $\pm$1.51 80.52 $\pm$1.18 76.35 $\pm$1.57 83.40 $\pm$3.00 81.31 $\pm$2.88
GPF  55.04 $\pm$1.82 54.20 $\pm$2.03 60.48 $\pm$1.11 55.57 $\pm$0.97 64.96 $\pm$6.77 60.63 $\pm$5.17 73.32 $\pm$2.56 68.53 $\pm$2.21 83.08 $\pm$0.68 80.60 $\pm$1.04
GPF-plus  56.36 $\pm$1.46 55.84 $\pm$1.33 61.36 $\pm$0.70 56.50 $\pm$0.44 72.32 $\pm$4.17 67.52 $\pm$3.51 80.56 $\pm$1.97 76.69 $\pm$1.87 85.08 $\pm$1.22 82.26 $\pm$0.94
SUPTsoft  55.24 $\pm$1.01 54.69 $\pm$1.18 59.60 $\pm$1.97 54.88 $\pm$1.45 62.40 $\pm$7.67 58.76 $\pm$6.17 77.28 $\pm$4.18 72.63 $\pm$4.86 84.52 $\pm$1.55 82.22 $\pm$1.36
SUPThard 56.96 $\pm$1.28 55.86 $\pm$1.28 61.28 $\pm$1.84 56.81 $\pm$1.45 71.64 $\pm$5.98 65.20 $\pm$5.70 77.96 $\pm$1.61 73.60 $\pm$1.72 85.24 $\pm$0.81 82.92 $\pm$0.56
GPPT  45.76 $\pm$0.34 46.05 $\pm$0.30 57.00 $\pm$0.55 51.03 $\pm$0.68 71.36 $\pm$0.70 64.53 $\pm$0.84 75.80 $\pm$0.42 69.97 $\pm$0.57 82.48 $\pm$0.74 80.65 $\pm$0.71
RELIEF 56.76 $\pm$0.23 55.94 $\pm$0.25 65.16 $\pm$1.21 58.55 $\pm$1.12 72.36 $\pm$1.05 66.71 $\pm$1.10 77.68 $\pm$0.93 72.53 $\pm$1.26 85.16 $\pm$0.89 84.07 $\pm$0.69
ContraEdge FT 60.60 $\pm$1.89 59.82 $\pm$2.01 60.88 $\pm$2.79 56.90 $\pm$2.20 72.20 $\pm$1.02 67.44 $\pm$0.81 80.64 $\pm$1.23 76.02 $\pm$1.75 82.24 $\pm$1.92 79.71 $\pm$1.79
GPF  59.12 $\pm$1.95 58.70 $\pm$1.92 63.24 $\pm$3.09 55.00 $\pm$2.54 71.24 $\pm$1.98 66.35 $\pm$2.84 80.68 $\pm$1.48 76.29 $\pm$1.60 83.64 $\pm$1.30 81.33 $\pm$1.27
GPF-plus 60.60 $\pm$1.08 60.63 $\pm$0.87 62.44 $\pm$1.86 56.81 $\pm$0.55 73.64 $\pm$3.73 69.95 $\pm$3.67 80.68 $\pm$0.65 76.93 $\pm$0.92 86.20 $\pm$1.15 83.57 $\pm$1.19
SUPTsoft  60.28 $\pm$1.32 60.05 $\pm$1.08 63.16 $\pm$2.35 55.95 $\pm$1.69 73.52 $\pm$3.81 69.70 $\pm$3.31 80.84 $\pm$0.85 75.30 $\pm$0.81 85.36 $\pm$1.75 82.47 $\pm$1.94
SUPThard  58.92 $\pm$0.86 58.75 $\pm$0.82 64.32 $\pm$2.38 55.96 $\pm$2.20 73.56 $\pm$1.97 69.48 $\pm$2.18 80.52 $\pm$0.95 76.69 $\pm$0.87 86.28 $\pm$0.55 83.92 $\pm$0.34
GPrompt  57.20 $\pm$2.13 53.98 $\pm$2.82 59.52 $\pm$0.65 55.46 $\pm$0.70 70.00 $\pm$0.62 66.17 $\pm$0.51 67.44 $\pm$0.80 64.81 $\pm$1.42 74.08 $\pm$2.55 71.26 $\pm$2.66
RELIEF 61.20 $\pm$1.20 60.89 $\pm$0.91 66.28 $\pm$0.27 58.01 $\pm$0.32 74.96 $\pm$1.72 70.22 $\pm$1.70 82.04 $\pm$0.83 78.03 $\pm$0.77 86.48 $\pm$0.90 83.88 $\pm$0.62
A.4. Experimental Settings
Data splitting and preprocessing.
We construct 10-shot node classification experiments by randomly sampling 10 nodes per class, with additional 500 nodes each for validation and testing. Considering the average degree of different datasets, we set varying $k$ for generating k-hop subgraphs, specifically 4, 3, 2, 2, 2 for Cora, Citeseer, Pubmed, Computers and Photo, respectively, to ensure an appropriate number of nodes for induced subgraphs.
Implementation.
We employ a 2-layer GIN (Xu et al., 2018) as the fundamental GNN model, which is widely utilized in previous works (Kipf and Welling, 2017; Shchur et al., 2018; Sun et al., 2022), with a hidden dimension of 128. While more advanced subgraph GNN techniques might yield better performance, as noted in (Fang et al., 2024), we opt to use the basic GIN for processing prompted subgraphs to ensure a fair comparison with GPPT and GraphPrompt, as these two methods do not incorporate subgraph-related designs. To achieve more compact node representations, we use Singular Value Decomposition (SVD) to reduce the initial features of each dataset to 100 dimensions. Cross-entropy loss is used as the reward function. The policy network and the projection head are independently updated using two Adam optimizers. Other hyper-parameters are listed in Table 6. The RL hyper-parameter settings are partly referenced from (Jiang et al., 2021), aligning with its established terminology.
Table 6. Hyper-parameter settings for graph and node level tasks.
Hyper-parameter graph-level node-level
General Hyper-parameters
# Training epochs {50, 100} {50, 100, 300}
Graph loader batch size {8, 16, 32, 64}
# GIN layers 5 2
GIN hidden dimension 300 128
Dropout ratio 0.5
Actors learning rate (lr) 5e-4
# Sub-policies $l$ 3
Discrete actors penalty $\alpha_{d}$1e0, 1e1, 1e3, 1e5}
Continuous actors penalty $\alpha_{c}$ 0.1
Critic lr 5e-4
Maximum prompt scale $z_{\text{max}}$0.1, 0.5, 1.0} {0.05, 0.1, 0.5}
Policy network weight decay 1e-5
Policy network lr linear decay {True, False}
Projection head lr {5e-4, 1e-3, 1.5e-3}
# Projection head layers {1, 2, 3}
# Projection head updates $q$1, 2, 3}
Projection head weight decay 0 5e-4
Projection head lr linear decay {True, False}
RL Hyper-parameters
Discount factor $\gamma$ 0.99
GAE $\lambda$ 0.95
# PPO epochs 10
PPO mini-batch size {32, 64, 128, 256, 512}
PPO clip range 0.2
Return normalization True
Entropy bonus coefficient $\beta$ 0.01
Critic loss coefficient 0.5
A.5. Performance Results
To evaluate the prompt tuning methods, we conducted five rounds of experiments with different random seeds, assessing performance using accuracy and macro F1-score. As illustrated in Table 5, RELIEF consistently achieves top-tier performance across majority of tasks, frequently ranking either first or as the runner-up, underscoring its applicability and effectiveness. On average, RELIEF outperforms fine-tuning by 2.05% and 1.43% on accuracy and macro f1-score, respectively. Additionally, RELIEF surpasses the second-best method, GPF-plus, by 0.89% in accuracy.
In contrast, GPPT and GraphPrompt exhibit inferior performance in most tasks. A plausible explanation is that their prediction results rely on class prototypes, which are challenging to accurately construct in few-shot scenarios with limited labeled data.
Figure 5. The testing accuracy of RELIEF and GPF-plus with and without feature prompts.
[…]
Figure 5 Fully described in the text.
A.6. Case Study
One notable observation, as demonstrated in Table 5, is that RELIEF’s performance lags behind its usual advantage in other tasks on the Computers dataset with GIN pre-trained by MaskedEdge. We thus conduct a case study to investigate the underlying causes.
Despite employing various widely-used training techniques – such as dynamically reducing the learning rate, altering batch sizes, applying weight decay in the Adam optimizer, and adjusting the dropout ratio – the self-supervised pre-training loss of the GIN model in this case still fails to decrease steadily. This indicates inadequate pre-training of the GNN model.
To investigate the discrepancy between GPF-plus (which performs best under this scenario) and RELIEF, we examine testing accuracy curves with and without incorporating prompts. As depicted in Figure 5, for GPF-plus, the testing accuracy without prompts rises at early epochs, reaching a maximum of 20.0, but then almost monotonically decreases to around 7.0, worse than random guessing in the ten-class classification task on the Computers dataset. Conversely, for RELIEF, although there is a downward trend in the first 50 epochs, with a minimum accuracy of 25.8, the accuracy then rapidly increases to about 75.1 and gradually converges.
We attribute the discrepancy in testing accuracy to different extents of disturbance caused by feature prompts, supported by two aspects. Firstly, we compare the APM (Average Prompt Magnitude) of GPF-plus and RELIEF, which are 0.194 and 0.053, respectively, against the original feature scale of 0.502, computed by the mean of L1-norm of each initial node feature averaged on dimension. Additionally, GPF-plus prompts on all nodes while RELIEF has a PCR (Prompt Coverage Ratio) of 0.61. Secondly, we apply linear probing (Kumar et al., 2022) to this case by solely adjusting the projection head, which is widely used in NLP domain to evaluate the expressiveness of the pre-trained model. The final testing accuracy of linear probing is 68.6, as indicated by a horizontal dotted gray line in Figure 5. When discarding prompts, GPF-plus’s accuracy plummets from 68.6 to 7.0, implying it overrides the initial features with overwhelming prompts, resulting in the pre-trained representations (obtained by directly feeding the original node features into the pre-trained GNN) becoming ineffective. In contrast, RELIEF maintains pre-trained knowledge and further generalizes it, leading to an accuracy increase from 68.6 to 77.6.
This raises the question: Why does GPF-plus excel while RELIEF struggles to show its effectiveness? With an inadequately pre-trained GNN model, which neither provides differentiated representations based on initial features nor adjusts itself during prompt tuning, GPF-plus relies on larger feature prompts to create differences in the input space. As a result, it generates discriminative representations better aligned with training labels via the collaboration between prompts and the projection head.
However, this scenario poses challenges for our RL-based RELIEF. As discussed in Section 3.2, the essence of RL is to map states to actions or to state-value functions. Actors are guided by state-values to explore and update, with state-values computed based on rewards. Thus, capturing the relationships between states and rewards is crucial for a powerful agent. In our scenario, states are constructed on node representations, and rewards are derived from the loss function. However, with the GNN model’s insufficient capacity, it may provide similar representations for samples with different labels and distinct representations for samples of the same class. Consequently, a trajectory (composed of transitions sampled on a single graph during exploration) effective on one graph may fail on another graph with similar representations but different labels. This leads to contradictory transitions, distracting the critic from accurately estimating state-value functions, thereby increasing the difficulty of obtaining the optimal policy.
In summary, GPF-plus tends to overshadow the original input to build a new mapping when faced with an inadequately pre-trained GNN model, while RELIEF performs well by interacting in an environment with stable reward dynamics provided by adequately pre-trained models. Therefore, RELIEF and GPF-plus can excel in complementary settings, where well-pre-trained models are more likely to be encountered and are preferable.
Appendix B More Information on Graph Classification
For datasets and pre-training strategies used in graph classification, we adhere to the training procedures established by pioneering researchers (Hu et al., 2019), with pre-trained GIN models available on their GitHub repository  [^1] https://github.com/snap-stanford/pretrain-gnns/tree/master/chem/model_gin.
B.1. Details of Datasets
Chemical datasets (Hu et al., 2019) are used to pre-train GNN models, consisting of both unlabeled and labeled molecules to facilitate comprehensive model training. Specifically, 2 million unlabeled molecules from the ZINC15 database are utilized for node-level self-supervised pre-training. For graph-level multi-task supervised pre-training, a preprocessed subset of the ChEMBL database including 456k labeled molecules is used, offering a diverse chemical space that enhances the model’s ability to generalize across various biochemical tasks.
For downstream graph classification tasks, 8 binary classification datasets for molecular property prediction about biophysics and physiology (Wu et al., 2018) are employed. The statistics of these downstream datasets are summarized in Table 7.
Table 7. Statistics of the downstream datasets used in our graph classification experiments.
Dataset # Molecules # Tasks   Dataset # Molecules # Tasks
BBBP 2039 1   ClinTox 1478 2
Tox21 7831 12   MUV 93087 17
ToxCast 8575 617   HIV 41127 1
SIDER 1427 27   BACE 1513 1
B.2. Details of Pre-training Strategies
Four pre-training strategies are used to assess the universal applicability of prompting approaches.
Deep Graph Infomax (Infomax) (Veličković et al., 2018) aims to derive expressive representations for graphs or nodes by maximizing the mutual information between the overall graph-level representations and the substructure-level representations at various granularities.
Attribute Masking (AttrMasking) (Hu et al., 2019) involves obscuring certain node or edge attributes. GNN models are then tasked with predicting these masked attributes based on the surrounding structural information.
Context Prediction (ContextPred) (Hu et al., 2019) uses subgraphs to predict their surrounding graph structures, which is designed to map nodes that appear in similar structural contexts to nearby embeddings, thereby capturing the context in which nodes are found.
Graph Contrastive Learning (GCL) (You et al., 2020) focuses on embedding augmented versions of the same anchor closely together (positive samples) while pushing the embeddings of different samples (negatives) apart.
The GNN models are first pre-trained using one of the above self-supervised methods and follow by the graph-level multi-task supervised pre-training.
B.3. Experimental Settings
We divide each dataset into training, validation, and testing sets with ratios of 80%, 10%, and 10%, respectively. Molecules are divided based on their scaffolds, i.e., molecular graph substructures (Ramsundar et al., 2019), and the clusters are recombined by prioritizing the most frequently occurring scaffolds in the training set. This produces validation and testing sets containing structurally different molecules (Hu et al., 2019), allowing us to evaluate the model’s out-of-distribution generalization.
To create 50-shot scenarios, we randomly sampled 50 graphs from the training set, contrasting with the original 1.1k to 74k training samples. This procedure introduced randomness, differing from previous works (Fang et al., 2024; Lee et al., 2024) that selected training data in descending order of scaffold occurrence. Given the limited budget of only 50 samples, the prior method resulted in training data with very few scaffold types, exacerbating overfitting. Consequently, the average ROC-AUC of fine-tuning across tasks under the 50-shot setting reported in previous works (Fang et al., 2024; Lee et al., 2024) was 55.40±2.13%, slightly above a random guess for binary classification tasks, thus potentially reducing the performance discrepancy between each prompting method. In contrast, our randomness brings about a dataset with more molecular diversity and also prevents scaffold overlap between training, validation and test sets. As a result, the average ROC-AUC of fine-tuning under our data splitting method is 60.92% with a smaller standard deviation of 1.46%.
We employ a 5-layer GIN with hidden dimensions of 300 as the base model. The policy network and the projection head are independently updated using two Adam optimizers. We assess downstream task performance using the validation early stopping protocol, reporting the test ROC-AUC at the best validation epoch. Additionally, considering that 50-shot learning may be insufficient for some datasets to ensure stable performance at the onset of training, we determine the best validation epoch after 20 epochs. The complete hyper-parameters for graph classification are detailed listed in Table 6.
B.4. Tuning Process Analysis
We analyze RELIEF’s tuning process on BACE using a GIN pre-trained with Infomax. In Figure 3, “Reward” denotes the cumulative reward for each graph, and ”Average Reward” is the mean “Reward” over a graph set.
Figure 3(a) shows that the ROC-AUC scores for the training, validation, and testing sets consistently increase during tuning, with validation performance converging around 50 epochs. Figure 3(b) illustrates the average reward induced by RELIEF’s prompting policy. Since scaffold splitting results in similar data distributions for the training and validation sets, their reward curves are also similar, increasing during the first 35 epochs and then gently decreasing. This decrease is due to the penalty aimed at improving policy generalization. As sub-policies diverge and produce distinct actions, the penalty – which measures the divergence between each sub-policy and the joint policy – increases. This prevents the joint policy from adhering to local contexts, thus enhancing generalization, though it leads to a rewards drop. Consequently, the testing set’s average reward rises steadily, indicating good generalization.
Figure 3(c) and (d) depict the reward distributions for the training and testing sets from 20 to 50 epochs, using the same random seed. The white vertical lines indicate average rewards at different epochs, showing a similar trend to Figure 3(b). In the early stages, such as at 20 epochs, rewards hover near 0. This resembles attaching random feature prompts to the original feature space, akin to data augmentation, which contributes to a more robust projection head. As tuning progresses, the policy network learns to map states to actions more discriminatively, resulting in broader reward distributions with an almost monotonically increasing mean for the testing set.
Figure 6. Effectiveness of discrete and continuous policies.
[…]
Figure 6 Fully described in the text.
Appendix C Ablation Study
Considering the core of RELIEF lies in its two actors for incorporating feature prompts, we deactivated either one to create two variants. Specifically, “random_d” replaces the discrete policy by randomly sampling a valid node to prompt at each step, while “random_c” replaces the continuous policy with a real-valued vector sampled from a random Gaussian distribution. Additionally, we include “linear probing” (Kumar et al., 2022), which discards the prompting components and adjusts only the projection head.
Figure 6 shows the comparisons of RELIEF with its variants on three graph-level and two node-level tasks. We find that deactivating either actor results in an obvious performance drop due to the disruption of their coordination. Particularly, the discrete actor plays a crucial role in RELIEF. Intuitively, a random prompt context assigned to a target node can be rectified by prompting on the node again in later steps, given the premise of a powerful discrete actor for locating this node. However, with a random node selection policy, a suboptimal feature prompt is difficult to mitigate by prompting on arbitrary nodes, and is further exacerbated by the message-passing mechanism, which spreads chaotic prompts throughout the graph. In addition, ”linear probing”, which only trains the projection head, exhibits significantly inferior performance, clearly demonstrating the contribution of the inserted feature prompts.
Appendix D Parameter Analysis
Figure 7. Parameter sensitivity analysis.
[…]
Figure 7 fully described in the text.
We analyze the critical hyper-parameters in RELIEF, including the maximum feature prompt scale per step $z_{\text{max}}$, the number of sub-policies $l$, and the two penalty factors $\alpha_{d}$ and $\alpha_{c}$ related to policy generalization. Figure 7 illustrates the ROC-AUC on ClinTox and accuracy on Computers, representing graph and node-level tasks, with each hyper-parameter varied over its legal range of values. The observations are as follows:
From Figure 7, we have the following observations. RELIEF performs consistently well across a broad range of hyper-parameter values, demonstrating its robustness. Specifically, when $z_{\text{max}}$ is too large, the excessive magnitude of prompts introduces unnecessary disturbance, leading to performance degradation. Conversely, setting $z_{\text{max}}$ too small weakens the impact of feature prompting, making it more similar to linear probing. By setting $l$ to 1, which discards the policy generalization technique, results in significant performance decline. While increasing $l$ may improve performance, it also requires maintaining more pairs of actors, leading to computational overhead. For $\alpha_{d}$ and $\alpha_{c}$, excessively large values may result in overly punishment, making the policies too conservative and hindering performance improvements.
Appendix E Introduction to PPO and H-PPO
Under our feature prompting scenario, an MDP is denoted as a tuple $\left\langle\mathcal{S},\mathcal{A}_{d},\mathcal{A}_{c},P,r,\gamma\right\rangle$, where $\mathcal{S}$ is the state space, and $(a,z)$, with $a\in\mathcal{A}_{d}$ and $z\in\mathcal{A}_{c}$, represents a hybrid action, $\gamma\in\left[0,1\right)$ is the discount factor. At each step, the agent receives a reward $r(s,a,z,s^{\prime})\in\mathbb{R}$ for performing an action $(a,z)$ in state $s$ and transitions to the next state $s^{\prime}$ with probability $P(s^{\prime}|s,a,z)$. We wish to find the policies $\pi_{d}(a|s)$ and $\pi_{c}(z|s,a)$, which determine a discrete action and a corresponding continuous action, respectively, for each state, to maximize the expected sum of discounted rewards, i.e., the return $R=\sum^{T}_{t=0}\gamma^{t}r(s_{t},a_{t},z_{t},s_{t+1})$, where $T$ is the maximum time step in a finite-horizon setting.
E.1. PPO
To maximize the return, policy gradient methods are typically employed, directly improving a stochastic policy $\pi_{\Phi}$ using gradient ascent. A general policy gradient method aims to maximize the following objective:
$\nabla_{\Phi}J(\Phi)=\mathbb{E}_{(s,a,s^{\prime})\sim\pi}\left[\sum_{t=0}^{T}\nabla_{\Phi}\log\pi_{\Phi}(a_{t}|s_{t})\mathbf{M}_{t}\right]$
where $\mathbf{M}_{t}$ represents any value-oriented metric that guides the action distribution output by the policy towards high-value regions. A widely used choice for $\mathbf{M}_{t}$ is the advantage function $A(s_{t},a_{t})$, which can be estimated using Generalized Advantage Estimation (GAE) (Schulman et al., 2015).
However, vanilla policy gradient methods do not constrain the extent of policy changes, which can lead to large updates and potential collapse in policy performance. To mitigate the risk of aggressive updates, PPO (Schulman et al., 2017) constrains the new policy to remain close to the old one by incorporating a clipped policy ratio $r(\Phi)$ within the objective, forming the surrogate objective $\mathcal{L}^{\text{surr}}$, defined as:
$\mathcal{L}^{\text{surr}}(\Phi)&=\mathbb{E}_{\pi}\left[\min\left(r(\Phi)A(s,a),\text{clip}\left(r(\Phi),1-\epsilon,1+\epsilon\right)A(s,a)\right)\right],\end{aligned}\\[5.0pt] \begin{aligned} r(\Phi)&=\frac{\pi_{\Phi}(a|s)}{\pi_{\Phi\text{old}}(a|s)}\end{aligned}\end{array}$
To encourage adequate exploration during training, an entropy regularizer is applied, promoting the policy to try diverse actions. This leads to the entropy-regularized PPO surrogate objective $\mathcal{L}^{\text{PPO}}$:
(10) $\mathcal{L}^{\text{PPO}}(\Phi)=\mathcal{L}^{\text{surr}}(\Phi)+\mathbb{E}_{\pi}[\beta\mathcal{H}(\pi_{\Phi}(a|s))]$
where $\beta$ is the entropy bonus coefficient. While the actors are updated using $\mathcal{L}^{\text{PPO}}$ to achieve higher returns, the critic is responsible for providing accurate state-value estimations, which are then used to compute the advantage function via GAE. The loss function for the critic, $\mathcal{L}^{\text{Critic}}$, commonly takes the form of Mean Squared Error (MSE), is expressed as:
(11) $\mathcal{L}^{\text{Critic}}(\Psi)=\mathbb{E}_{\pi}\left[\left(V_{\Psi}(s_{t})-R_{t}\right)^{2}\right],\quad R_{t}=\sum_{l=0}^{\infty}\gamma^{l}r_{t+l}$
Notably, PPO can be applied to either discrete or continuous action spaces, but not simultaneously. For further implementation details, please refer to (Huang et al., 2022) and our code.
E.2. Hybrid Action Space and H-PPO
A hybrid action space commonly exhibits a hierarchical structure: a high-level discrete action is chosen first, followed by a low-level parameter, treated as a continuous action associated with the discrete action. H-PPO (Fan et al., 2019) is based on the actor-critic framework but enhances it by employing two parallel actors – one for selecting discrete actions and another for determining continuous actions – while a single critic estimates the state-value function.
The training schema of H-PPO is based on PPO, with the main difference being that the two actors are updated separately. Specifically, after sampling a batch of transitions $(s,a,z,r,s^{\prime})$, the discrete actor is optimized by maximizing Eq. (10), denoted as $\mathcal{L}^{\text{PPO}}(\Phi_{d})$, using $(s,a,r,s^{\prime})$, whereas the continuous actor is optimized in the same form, denoted as $\mathcal{L}^{\text{PPO}}(\Phi_{c})$, using $(s,z,r,s^{\prime})$. Meanwhile, the critic is updated by minimizing Eq. (11), comparing its estimated state-values with the true return values based on $(s,r,s^{\prime})$.
During evaluation, a state is first sent to the discrete actor and outputs an action probability distribution, from which a discrete action is sampled. The same state is subsequently sent to the continuous actor and outputs continuous parameters for each discrete action. The specific parameter corresponding to the selected discrete action is fetched, and a continuous action is sampled from a Gaussian distribution based on the selected parameter as the mean.
Appendix F Pseudo-Code
RELIEF comprises two trainable modules: the policy network $\Pi_{\omega}$ and the projection head $g_{\phi}$. The policy network $\Pi_{\omega}$ includes $l$ discrete and continuous actors, respectively, and a single critic. The actors represent sub-policies ${\pi_{\Phi_{d,1}},\dots,\pi_{\Phi_{d,l}}}$ and ${\pi_{\Phi_{c,1}},\dots,\pi_{\Phi_{c,l}}}$, while the critic represents the state-value function $V_{\Psi}(s)$. Thus, the training parameters are ${\Phi_{d,1},\dots,\Phi_{d,l},\Phi_{c,1},\dots,\Phi_{c,l},\Psi,\phi}$. For simplicity, we use $\omega$ to denote the parameters of all the actors and the critic. The pseudo-code for training RELIEF, which alternates between training the policy network and the projection head, is presented in Algorithm 1. The operations in line 7 and line 22 can be batchified by consuming a batch of graphs.
[…]
Note that at the evaluation stage, the joint policies $\pi_{d,\,J}$ and $\pi_{c,\,J}$ are utilized to attach feature prompts, similar to the projection head training, so it is not further described here.
Algorithm 1: Training RELIEF
1:The frozen pre-trained GNN model $f_{\theta}$, training graphs $\mathcal{D}=\{(\mathcal{G}_{1},y_{1}),\dots,(\mathcal{G}_{m},y_{m})\}$ corresponding to node numbers $\{n_{1},\dots,n_{m}\}$, the number of sub-policies $l$, the maximum number of nodes $N$, the number of times the projection head is trained per epoch $q$, and the maximum training epoch $E$.
2:Optimal policy network $\Pi_{\omega}$ and projection head $g_{\phi}$.
3:Initialize parameters $\omega$ and $\phi$;
4:Bootstrap sampling $l$ training graph sets $\{\mathcal{D}_{1},\dots,\mathcal{D}_{l}\}$ from $\mathcal{D}$ for $l$ actors;
5:for $epoch=1$ to $E$ do
6:     for $i=1$ to $l$ do $\triangleright$ Policy network training
7:         for each $(\mathcal{G}_{j},y_{j})$ in $\mathcal{D}_{i}$ do
8:              Compute initial loss $\mathcal{L}_{0}=\mathcal{L}(g_{\phi}(f_{\theta}(\mathcal{G}_{j})),y_{j})$;
9:              Initialize a prompt feature matrix $\mathbf{P}_{0}$ with zeros;
10:              Initialize an experience replay buffer $\mathcal{B}$.
11:              for $t=1$ to $n_{i}$ do
12:                  Construct state $s_{t}$ by Eq. (1);
13:                  Determine a feature prompt $p_{t}^{a,z}$ by sub-policies $\pi_{\Phi_{d,i}}$ and $\pi_{\Phi_{c,i}}$ and adds to $\mathbf{P}_{t-1}$;
14:                  Compute instant reward $r_{t}$ by Eq. (2);
15:                  Save the transition $(s_{t},a_{t},z_{t},r_{t},s_{t+1})$ to $\mathcal{B}$;
16:              end for
17:              Update discrete actor $\pi_{\Phi_{d,i}}$, continuous actor $\pi_{\Phi_{c,i}}$ and critic $V_{\Psi}(s)$ with $\mathcal{B}$ by Eq. (4), Eq. (6) and $\mathcal{L}^{\text{Critic}}$;
18:         end for each
19:     end for
20:
21:     Initialize an empty list $\mathcal{D}^{\ast}$; $\triangleright$ Projection head training
22:     for each $(\mathcal{G}_{k},y_{k})$ in $\mathcal{D}$ do
23:         Initialize a prompt feature matrix $\mathbf{P}_{0}$ with zeros;
24:         for $t=1$ to $n_{k}$ do
25:              Determine a feature prompt $p_{t}^{a,z}$ by joint policies $\pi_{d,\,J}$ [Eq. (5)] and $\pi_{c,\,J}$ [Eq. (7)] and adds to $\mathbf{P}_{t-1}$;
26:         end for
27:         Save the prompted graph $\mathcal{G}_{k}^{\ast}$ to $\mathcal{D}^{\ast}$;
28:     end for each
29:     for $times=1$ to $q$ do
30:         Update the projection head with $\mathcal{D}^{\ast}$ by Eq. (3);
31:     end for
32:end for
Appendix G Complexity Analysis
We analyze the time complexity of training and evaluation RELIEF for one epoch. The GNN model is a $L_{g}$-layer GIN with a hidden size $d_{g}$. By neglecting the dimension discrepancy between actors and the critic, we assume that the policy network is equipped with $(2l+1)$ $L_{p}$-layer MLPs with a hidden size $d_{p}$, where $l$ is the number of sub-policies. The projection head is a $L_{h}$-layer MLP with a hidden size $d_{h}$. In addition, $N$, $E$ are the maximum number of nodes and edges of $m$ and $m^{\prime}$ graphs in training and evaluation sets, respectively, and $D$ is the initial dimension of node features.
G.1. Detailed Complexity
The training process involves agent sampling, policy network training and projection head training.
Agent sampling.
As we discuss in Section 3.2, the trainable components of actors and the critic are essentially MLPs. Starting with a pair of discrete and continuous actors, to determine prompt features at each step for one graph, the state with dimension $d_{g}$ is passed to the two actors, with a complexity of $2O(d_{g}d_{p}+L_{p}d_{p}^{2})$. The prompted graph is then sent to the GIN to obtain node representations, with a complexity of $O(L_{g}(Ed_{g}+NDd_{g}))$. Using mean pooling, the node representations are transformed into a graph representation, with a complexity of $O(Nd_{g})$. Subsequently, by feeding these representations into the projection head, it outputs predictions with a complexity of $O(d_{g}d_{h}+L_{h}d_{h}^{2})$, and computes the loss with $O(1)$. This process repeats $N$ times to collect $N$ transitions. Consequently, by sequentially performing sampling for each pair of actors and processing all $m$ graphs in the training set, the total time complexity is:
(12) $\begin{split}mN&\cdot[2O(d_{g}d_{p}+L_{p}d_{p}^{2})\\ &+O(L_{g}(Ed_{g}+NDd_{g})+Nd_{g})+O(d_{g}d_{h}+L_{h}d_{h}^{2})+O(1)]\end{split}$
Policy network training.
A total of $mN$ transitions are used for sequentially updating $l$ pairs of actors, along with the critic. Therefore, the complexity of policy network training, which essentially updates MLPs, is:
(13) $3mN\cdot O(d_{g}d_{p}+L_{p}d_{p}^{2})$
Projection head training. This process is similar to agent sampling, except that it uses joint discrete and continuous policies to incorporate prompts. Then, the projection head is updated $q$ times based on the prediction loss. Thus, the time complexity of projection head training is:
(14) $\begin{split}&mNl\cdot[2O(d_{g}d_{p}+L_{p}d_{p}^{2})+O(L_{g}(Ed_{g}+NDd_{g})+Nd_{g})]\\ &+mq\cdot[O(d_{g}d_{h}+L_{h}d_{h}^{2})+O(1)]\end{split}$
To sum up, the time complexity of training one epoch is the sum of Eq. (12), Eq. (13) and Eq. (LABEL:eq:complex_projection_training).
The time complexity during the evaluation stage is similar to that of projection head training, since it also involves inserting prompts using joint policies, but making predictions with the projection head only once, rather than $q$ times. Therefore, the time complexity of evaluation can be derived by replacing $q$ with 1 and $m$ with $m^{\prime}$ in Eq. (LABEL:eq:complex_projection_training).
G.2. Simplified Complexity
Given the cumbersome expression of time complexity, we aim to simplify and approximate it for clearer insight. Specifically, we unify all the hidden sizes $d_{g}$, $d_{p}$ and $d_{h}$ as $d$, and ignore small variables as constants, such as the number of layers $L_{(\cdot)}$ and training times $q$ for the projection head. Then, the simplified time complexity of training one epoch is:
$m\cdot O(N^{2}Dd+N^{2}d+Nd^{2}+NED+d^{2}+N+1)$
and the simplified time complexity of evaluation is:
$m^{\prime}\cdot O(N^{2}Dd+N^{2}d+Nd^{2}+NED+d^{2}+1)$
Since the dimension $D$ and $d$ are fixed given specific application scenarios, the main bottleneck of the framework lies in $O(mN^{2}Dd)$ and $O(mNED)$, indicating that the computation time will noticeably increase with the number of large graphs scaling up. Fortunately, the results in Section 5.2 demonstrate that RELIEF is data efficient, meaning that with a few samples, i.e., small $m$, RELIEF can already achieve desired performance. Furthermore, more efficient exploration methods can be applied to RELIEF to decrease the sampling cost, which is left for future work.